The 2014 balance sheet of Marias Tennis Shop, Inc., showed long-term debt of $6.5 million, and the 2015 balance sheet showed long-term debt of $5.5 million. The 2015 income statement showed an interest expense of $49787. During 2015, Marias Tennis Shop, Inc. realized the following: Cash flow to creditors $16878 Cash flow to stockholders $71136 Suppose you also know that the firms net capital spending for 2015 was $2.2 million, and that the firm reduced its net working capital investment by $85565. What was the firms 2015 operating cash flow, or OCF?
Answer #1

Cash flow from assets is = Cash flow to creditors + cash flow to stockholders equity

Cash flow from assets is = (16878+71136)

Cash flow from assets is = $ 88,014/.

Cash flow from assets = OCF - Change in working capital - Net capital spending

88014 = OCF - (-85565) - (2.2*1000000)

OCF is = 2200000 - 85565 - 88014

OCF is = $ 2,026,421/.